You are here

Hermes/Nestor

Hermes/Nestor System Status

Post date System Status: Update Notes
2017-06-19 - 15:40 PDT System to be defunded

System Defunded - Local User Access ONLY

Local UVIC jobs ONLY after June 1. Storage access for any user until July 15.

Refer to the Compute Canada Documentation Wiki WestGrid Migration Guide for migration instructions. 

Contact sysadmin@uvic.ca for local use policies.

2017-05-18 - 08:57 PDT Online

File system issues on Hermes/Nestor.

Update:

The issue is resolved. No job were affected during this incident.

Original

We are having issues with GPFS on Hermes/Nestor on one of our headnodes. We are working on the issue and we will update this notice accordingly.

Sorry for the inconvenience.

 

2017-05-17 - 14:42 PDT Conditions

File system issues on Hermes/Nestor.

We are having issues with GPFS on Hermes/Nestor on one of our headnodes. We are working on the issue and we will update this notice accordingly.

Sorry for the inconvenience.

 

2017-05-03 - 14:38 PDT Online

Scheduled Logical Network Change Thursday May 4, 2017 11:00 PDT

We will be making a logical network change between 11:00-11:30am PDT tomorrow and no outage is anticipated during this time.

2017-02-02 - 15:56 PST Online

Scheduled Logical Network Changes Hermes/Nestor/West Cloud

Scheduled work has been completed.

-----------

On Thursday Feb 2 between 13:00-16:00 PT there will be some logical network changes which may cause brief periods of packet loss during this maintenance window. We have tested the changes and users likely will not notice any issues during this window. 

2017-01-31 - 09:29 PST Conditions

Scheduled Logical Network Changes Hermes/Nestor/West Cloud

On Thursday Feb 2 between 13:00-16:00 PT there will be some logical network changes which may cause brief periods of packet loss during this maintenance window. We have tested the changes and users likely will not notice any issues during this window. 

2017-01-26 - 10:22 PST Online

Networking Issues Hermes/Nestor

Hermes/Nestor is now back on-line.

-------------------------------------------------------------------

Overnight we had a management networking issue with Hermes/Nestor which caused issues with GPFS and DNS, etc. All running jobs will need to be resubmitted and currently all nodes are off-line. We are working on resolving the issue and expect the cluster to be available in the next hour.

Pages